OpenStudy (anonymous):

A line passes through (3, –2) and (6, 2). Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

First find the slope between the two points. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug them in: \(m = \dfrac{2 + 2}{6 - 3}\) Can you simplify that? @Sward

OpenStudy (anonymous):

Just add 2 + 2 and subtract 6 - 3..

OpenStudy (anonymous):

so its 4/3